# What To Do (and Not to Do) with Causal Panel Analysis under Parallel Trends:

### v2

@@ 3.  HTE-Robust Estimators
In this section, we offer a brief overview and comparison of several recently introduced HTE-robust estimators. For a more comprehensive discussion, please refer to the SM.
@@
Another noteworthy difference lies in the settings in which these estimators are applicable: some function exclusively in settings with staggered treatment adoption, while others can accommodate scenarios with treatment reversals. Furthermore, these estimators diverge in terms of (1) how they select untreated observations as controls for treated units, (2) how they incorporate pre-treatment or exogenous covariates, and (3) the choice of the reference period. We discuss these details further below and in the ‘Survey of HTE-Robust Estimators’ and ‘Implementation Details’ sections of the SM.
@@
Research article Sun and Abraham (2021) Callaway and Sant’Anna (2021) de Chaisemartin and D’Haultfœuille (2020) IKW (2023) BJS (2023) LWX (2022)
Method known as interaction weighted csdid did_mulitple PanelMatch DIDM FEct
Key ID assumption Parallel trends Parallel trends Parallel trends Parallel trends Zero conditional mean Strict exogeneity
Finest estimand $\delta_{g,l}$ $\delta_{g,l}$ $\delta_{l}$ $\delta_{l}$ $\tau_{i,t}$ $\tau_{i,t}$
Comparison group Never-treated or last-treated Never-treated or not-yet-treated Matched stable group (not-yet-treated) Matched stable group (not-yet-treated) Imputed Counterfactual (not-yet-treated) Imputed Counterfactual (not-yet-treated)
Reference period(s) Period 0 An arbitrary pre-treatment period Untreated period Period 0 All pre-treatment periods All pre-treatment periods
Covariate adjustment Possible extension Outcome & propensity score modeling Possible extension Refined matched set & outcome modeling Outcome modeling Outcome modeling
@@
DID extensions are all built from local, $2\times 2$ DID estimates—hence our choice of terminology. The overarching strategy for these estimators is to estimate the DTE, $\delta_{l}$ (or $\delta_{g,l}$ for each cohort $g$ in the staggered setting), for each period since the most recent initiation of treatment, $l$, using one or more valid $2\times 2$ DIDs. By ‘valid,’ we mean that the DID includes (1) a pre-period and a post-period and (2) a treated group and a comparison group. The pre-period is such that all observations in both groups are in control, while the post-period is such that observations from the treated group are in treatment and those from the comparison group are in control. The choice of the comparison group is the primary distinction between estimators in this category. To obtain higher-level averages such as the ATT, we then average over our estimates of $\delta_{l}$ (or $\delta_{g,l}$), typically employing appropriate, convex weights.
@@
We cover two estimators in this category that are appropriate only for the staggered setting. Sun and Abraham (2021) propose an interaction-weighted (iw) estimator, which is a weighted average of $\delta_{g,l}$ estimates obtained from a TWFE regression with cohort dummies fully interacted with indicators of relative time to the onset of treatment. They demonstrate that each resulting estimate of $\delta_{g,l}$ can be characterized as a difference in the change in average outcome from a fixed pre-period $s<g$ to a post-period $l$ periods since $g$ between the treated cohort $g$ and the comparison cohorts in some set $\mathcal{C}$ [^9] This equivalence holds when the panel is balanced, i.e. there is no missing data. When there is missing data, the estimator from the saturated regression differs from one that directly estimates local DIDs, including the never-treated version of the next estimator we introduce, csdid. The authors recommend using $\mathcal{C}={\sup_{i}E_{i}}$, which is either the never-treated cohort or, if no such cohort exists, the last-treated cohort. By default, iw uses $l=0$ as the reference period and can accommodate covariates in the TWFE regression.
@@
Employing the same general approach, Callaway and Sant’Anna (2021) propose two estimators, one of which uses never-treated units ($\hat{\delta}^{CS-dr}_{nev}$) and the other not-yet-treated units ($\hat{\delta}^{CS-dr}_{ny}$) as the comparison group. We label these estimators collectively as csdid. Note that $\hat{\delta}^{CS-dr}_{nev}$ uses the same comparison group as iw when a never-treated cohort exists, whereas $\hat{\delta}^{CS-dr}_{ny}$ uses all untreated observations of not only never-treated units but also later adopters as controls for earlier adopters. Besides the choice of comparison cohort, csdid estimators differ from iw in that they allow users to condition on pre-treatment covariates using both an explicit outcome model and inverse probability weighting (IPW) simultaneously, with at least one needing to be correct for the estimator to be consistent. We will only implement $\hat{\delta}^{CS-dr}_{ny}$, as most models in our replication sample do not include additional covariates, rendering $\hat{\delta}^{CS-dr}_{nev}$ and $\hat{\delta}^{IW}$ numerically identical. While iw uses one period before the treatment’s onset as the reference period, csdid allows users to choose one or multiple pre-treatment periods as the reference
@@
 [^10] The “stacked” DID or regression is another related estimator sometimes used to address HTE concerns, but it is not HTE-robust. As described by Baker, Larcker and Wang (2022), it involves creating separate sub-datasets for each treated cohort by combining data from that cohort (around treatment adoption) and never-treated cohort data from the same periods. These cohort-specific datasets are then “stacked” to form a single dataset. An event study regression akin to Equation (3) is then run, using sub-dataset specific unit and time dummies. This method is akin to iw and the never-treated version of csdid without covariates; they all use the same data. However, stacked regression estimates a single DTE for a given relative period, rather than separate estimates for each cohort. Essentially, stacked DID is a DID extension using implicit, immutable weights selected by OLS, and is not HTE-robust because OLS does not assign cohort-proportional weights; the effects for lower variance (larger) cohorts are generally down-weighted relative to their size.
@@
In settings with treatment reversals, separate groups of researchers have converged on the same strategy for choosing a comparison group: matching treated and control observations that belong to units with identical treatment histories. IKW (2023) suggest one such estimator, PanelMatch, which begins by constructing a “matched set” for each pair $(i,t)$ such that unit $i$ transitions into treatment at time $t$. This matched set includes units that are both (1) not under treatment at time $t$ and (2) share the same treatment history as $i$ for a fixed number of periods leading up to the treatment onset. For each post-period $l$ periods since treatment adoption, it then estimates a local DID. It uses a fixed pre-period $s<t$ and the post-period $t-1+l$. The treatment “group” comprises solely of $(i,t)$, and the members of the matched set for $(i,t)$ that are still under control during the post-period serves as the comparison group. To obtain $\delta_{l}$ for a given $l$, it then averages over the local estimates from all treated $(i,t)$ such that $(i,t+l-1)$ is still under treatment (i.e., no treatment reversal has occurred yet). IKW (2023) propose incorporating covariates by “refining” matched sets and use $l=0$ as the reference period.
Using a similar strategy, de Chaisemartin and D’Haultfœuille (2020) propose a “multiple DID” estimator, DIDM. Importantly, they include local DIDs for units leaving treatment and not only those joining treatment. DIDM also only consider the case where we match on a single period and where $l=1$. Consequently, their target estimand is not the ATT but rather an average of the contemporaneous effects of “switching” (i.e., the effect of joining or the negative of the effect of leaving at the time of doing so). Interestingly, another DID extension can be seen as a special case of PanelMatch: In the staggered setting, the PanelMatch estimator aligns with the not-yet-treated version of csdid (without covariate adjustment). We delve into details on the connections between these three estimators in the ‘Survey of HTE-Robust Estimators’ section of the SM.
@@
All DID extensions are built using local, $2\times 2$ DIDs, and their assumptions reflect this. Specifically, they each rely on a form of the PTA—that is, the expected changes in potential outcomes under control from one period to the next are equal between the treated and the chosen comparison groups. In Table 1, we refer to all these assumptions collectively as the PTA. We defer readers to the SM for a fuller account of each method’s assumptions.
@@
Imputation estimators do not explicitly estimate local DIDs. Instead, they take the difference between the observed outcome and an imputed counterfactual outcome for each treated observation. The connection to the TWFE model is in the functional form assumption used to impute counterfactual outcomes. Specifically, an imputation estimator first fits a parametric model for the potential outcome under control $Y_{i,t}(0)$—in our case, this is Model (1)—using only control observations $\{(i,t):D_{i,t}=0\}$. It is also through this outcome model that one can adjust for covariates. Then, it imputes $Y_{i,t}(0)$ for all treated observations $\{(i,t):D_{i,t}=1\}$ using the estimated parameters. Finally, it estimates the ITE, $\tau_{i,t}$, for each treated observation $(i,t)$ by calculating the difference between its observed outcome $Y_{i,t}=Y_{i,t}(1)$ and its imputed counterfactual outcome $\hat{Y}_{i,t}(0)$. Inference for the estimated $\hat{\tau}_{i,t}$ is possible, although uncertainty estimates need to be adjusted to account for the presence of idiosyncratic errors (e.g., Bai and Ng, 2021). BJS (2023) and LWX (2022) each propose estimators in this category. Each group proposes a more general framework that nests many models, including TWFE. The latter also introduces several specific imputation estimators. One of these uses the TWFE model, and the authors refer to the resulting astimator as the fixed effect counterfactual estimator, or FEct.
@@
Compared to DID extensions, which typically use a single pre-period and, with the exception of csdid, only a subset of units under control at both the pre- and post-periods as a comparison group, imputation estimators use all available control observations to estimate treated counterfactuals. As such, for each unit, the reference period can be understood as (the average of) all pre-treatment periods. Intuitively, this approach should result in higher efficiency. In fact, BJS (2023) demonstrate that imputation estimators are the most efficient among all estimators under the condition of homoskedastic errors. Also in contrast to DID extensions, imputation estimators do not directly assume the PTA. Instead, they restrict the expectation of the error terms from a parametric TWFE model. In Table 2, this is denoted as “zero conditional mean” for BSJ (2023) or “strict exogeneity” for LWX (2022). Again, we refer readers to the SM for the formal statements of these assumptions. Collectively, these assumptions imply a form of conditional, baseline randomization of treatment assignment, which in turn implies the PTA (e.g., Blackwell and Glynn, 2018).
@@
Although DID extensions and imputation methods rely on slightly different identification assumptions, such as the PTA and specific constraints on the error terms, they usually lead to similar observable implications. Researchers commonly use the presence or absence of pretrends in the pre-treatment periods to judge the plausibility of the PTA. In the classic two-group, two-period setting, if there are data from additional pre-treatment periods, researchers can plot the time series of (average) outcomes of each group and visually inspect whether they indeed trend together. The intuition is that if the PTA holds and the outcome trends of the treated and control groups are indeed parallel when $Y(0)$’s are observed for all units, then it is plausible that the PTA also holds in the post-treatment periods, when $Y(0)$’s are no longer observable for units in the treatment group. Conversely, differential trends in the pre-treatment periods should make us suspicious of the PTA. In more complex settings or where we wish to control for observed confounders, we can substitute the outcome time series with estimates of the dynamic effects before the onset of treatment, $\delta_{l}$ for $l\leq 0$. If the PTA holds, then these pre-treatment effect estimates should be zero. We provide a more thorough discussion and examples of DTE plots, which are also known as “event study plots,” in the next section when we introduce our procedure.
@@ 4.  Data and Procedure
Next, we assess the robustness of empirical findings from causal panel analyses in political science and compare results obtained using the different methods we have discussed. We will explain our sample selection rules, describe standard practices in the field, and outline our reanalysis approach. Readers can find a more detailed explanation of our sample selection criteria and replication and reanalysis procedure in the “Sample and Replicability” and “Implementation Details” sections of the SM.
@@
Our replication sample comprises articles from three leading political science journals—the APSR, AJPS, and JOP—published over a recent six-year span from 2017 to 2022. We initially include all studies, including both long and short articles, that employ panel data analyses with a binary treatment as a crucial component of their causal argument, resulting in a total of 90 articles. After a careful review of each of these articles, we find that 52 articles use a TWFE model similar to Model (1). We then attempt to replicate the main results of these 52 articles and are successful in 37 cases (71.2%). A detailed explanation of how we select the “main model” is provided in the ‘Sample and Replicability’ section of the SM. Table 2 depicts the distribution of successful replications, along with reasons for replication failures, across the various journals.
@@
Table 3 presents an overview of the standard practices and settings in the articles that we successfully replicated. The vast majority of studies in our sample (92%) use DID designs to justify the use of the TWFE model, while the remaining studies advocate for the model’s ability to exploit “within” variations in the data. Out of the 37 articles, five (13%) employ a classic DID design, which includes two-group, two-period designs (three articles) and multi-period DID designs (two articles). 10 articles (27%) use a staggered (but not classic) DID design, while the remaining 22 articles (59.5%) fall into the “general” category, meaning they allow for treatment reversals. Except for four articles, all studies have a continuous outcome of interest. Most studies adopt cluster-robust SEs or panel-corrected SEs (Beck and Katz, 1995), while five articles apply bootstrap procedures for estimating uncertainties. A subset of authors explore alternative model specifications by adding LDVs (six articles), ULTs (11 articles) and higher-than-unit-level time trends (5 articles). Notably, 22 studies use some type of plot—either average outcomes over time, DTE/event-study plots, or both—to evaluate the plausibility of the PTA.
@@
Continuous 33 89.2% DTE/event-study plots 18 48.6%
@@
To begin, we aim to understand the research setting and data structure. We visualize the patterns of treatment and outcome variables using plots like Figure 3 (a) and (b). In this application, treatment reversals clearly take place. Some data is missing (due to redistricting), but the issue does not seem to be severe. We record important details such as the number of observations, units, and time periods, the type of variance estimators, and other specifics of the main model (not shown here). Next, we replicate the main finding, employing both the original variance estimator and a cluster-bootstrap procedure. We also use a bootstrap refinement procedure recommended by Cameron, Gelbach and Miller (2008) to conduct statistical testing; the results are broadly consistent with the ones based on cluster-bootstrapped SEs and CIs.
Figure 2: Visualizing Data from Grumbach and Sahn (2020)
(a) Treatment Status
(b) Outcome Variable
Note: Visualizing data from Grumbach and Sahn (2020), who investigate the effects of Asian candidates in congressional elections on the share of campaign contributions by Asian donors.
We then re-estimate the ATT and DTEs using estimators discussed in Section 3. For staggered adoption treatment cases, we apply six estimators: TWFE (with always treated units removed for easier comparisons with other estimators), PanelMatch, FEct, Stacked DID, iw, and csdid (not-yet-treated). For applications with treatment reversals like Grumbach and Sahn (2020), we use the first three estimators only. The comparison between the TWFE estimate and the other estimates sheds light on whether original findings are sensitive to the presence of HTE. Results from this example are shown in Figures 3(a)-(d). The similarity between estimates for ATTs in (a) suggests the robustness of the original finding to the choice of estimators. The DTE plots from HTE-robust estimators in (c)-(d) are broadly consistent with the DTE plot from TWFE in (b).
@@
Note: Reanalysis of data from Grumbach and Sahn (2020). Subfigures (a): Treatment effect or ATT estimates from multiple methods; subfigures (b)-(d): DTE plots using TWFE, PanelMatch, and FEct; Subfigure (e)-(f): results from the placebo test and test for carryover effects usingFEct—the blue points in (g) and red points in (h) represent the holdout periods in the respective tests. CIs for TWFE, PanelMatch, and FEct In all subfigures are produced by bootstrap percentile methods.

### v3

@@ 3.  HTE-Robust Estimators
In this section, we offer a brief overview and comparison of several recently introduced HTE-robust estimators. We use the term “HTE” to refer to ITE that are nonidentical, i.e. $\tau_{i,t}\neq\tau_{j,s}$ for some $i,j,s,t$, and “HTE-robust” to refer to estimators that identify some convex combination of ITE even in the presence of HTE. For a more comprehensive discussion on all estimators mentioned in this section, please refer to the SM.
@@
Another noteworthy difference lies in the settings in which these estimators are applicable: some function exclusively in settings with staggered treatment adoption, while others can accommodate scenarios with treatment reversals. [^10] It is possible to use some methods designed for the staggered setting even when there are reversals by using only a subset of the data, but this requires modifications to the original estimator and would identify a different effect than the ATT. In the latter setting, all estimators we discuss also require a “no carryover effects” assumption, i.e., the treatment effect does not persist after the current period, regardless of which estimator is used and whether it is HTE-robust. Furthermore, these estimators diverge in terms of (1) how they select untreated observations as controls for treated units, (2) how they incorporate pretreatment or exogenous covariates, and (3) the choice of the reference period. We discuss these details further below and in Sections A.1. and A.4. of the SM.
@@
Research article Sun and Abraham (2021) Callaway and Sant’Anna (2021) Baker, Larcker and Wang (2022) de Chaisemartin and D’Haultfœuille (2020) IKW (2023) BJS (2023) LWX (2022)
Method known as interaction weighted csdid StackedDID did_mulitple PanelMatch DIDimpute FEct
Key ID assumption Parallel trends Parallel trends Parallel trends Parallel trends Parallel trends Zero conditional mean Strict exogeneity
Finest estimand $\delta_{g,l}$ $\delta_{g,l}$ $\delta_{l}^{vw}$ $\delta_{l}$ $\delta_{l}$ $\tau_{i,t}$ $\tau_{i,t}$
Comparison group Never-treated or last-treated Never-treated or not-yet-treated Never-treated Matched stable group (not-yet-treated) Matched stable group (not-yet-treated) Imputed Counterfactual (not-yet-treated) Imputed Counterfactual (not-yet-treated)
Reference period(s) Period 0 An arbitrary pretreatment period Period 0 Untreated period Period 0 All pretreatment periods All pretreatment periods
Covariate adjustment Possible extension Outcome & propensity score modeling Outcome modeling Possible extension Refined matched set & outcome modeling Outcome modeling Outcome modeling
@@
DID extensions are all built from local, $2\times 2$ DID estimates—hence our choice of terminology. The overarching strategy for these estimators is to estimate the DTE, $\delta_{l}$ (or $\delta_{g,l}$ for each cohort $g$ in the staggered setting), for each period since the most recent initiation of treatment, $l$, using one or more valid $2\times 2$ DIDs. By “valid,” we mean that the DID includes (1) a pre-period and a post-period and (2) a treated group and a comparison group. The pre-period is such that all observations in both groups are in control, while the post-period is such that observations from the treated group are in treatment and those from the comparison group are in control. The choice of the comparison group is the primary distinction between estimators in this category. To obtain higher-level averages such as the ATT, we then average over our estimates of $\delta_{l}$ (or $\delta_{g,l}$), typically employing appropriate, convex weights.
@@
We cover three estimators in this category that are appropriate only for the staggered setting. Sun and Abraham (2021) propose an interaction-weighted (iw) estimator, which is a weighted average of $\delta_{g,l}$ estimates obtained from a TWFE regression with cohort dummies fully interacted with indicators of relative time to the onset of treatment. They demonstrate that each resulting estimate of $\delta_{g,l}$ can be characterized as a difference in the change in average outcome from a fixed pre-period $s<g$ to a post-period $l$ periods since $g$ between the treated cohort $g$ and the comparison cohort(s) in some set $\mathcal{C}$ [^11] This equivalence holds when the panel is balanced, i.e., there are no missing data. When there are missing data, the estimator from the saturated regression differs from one that directly estimates local DIDs, including the never-treated version of the next estimator we introduce, csdid. The authors recommend using $\mathcal{C}={\sup_{i}E_{i}}$, which is either the never-treated cohort or, if no such cohort exists, the last-treated cohort. By default, iw uses $l=0$ as the reference period and can accommodate covariates in the TWFE regression.
@@
Employing the same general approach, Callaway and Sant’Anna (2021) propose two estimators, one of which uses never-treated units ($\hat{\delta}^{CS-dr}_{nev}$) and the other not-yet-treated units ($\hat{\delta}^{CS-dr}_{ny}$) as the comparison group. We label these estimators collectively as csdid. Note that $\hat{\delta}^{CS-dr}_{nev}$ uses the same comparison group as iw when a never-treated cohort exists, whereas $\hat{\delta}^{CS-dr}_{ny}$ uses all untreated observations of not only never-treated units but also later adopters as controls for earlier adopters. Besides the choice of comparison cohort, csdid estimators differ from iw in that they allow users to condition on pretreatment covariates using both an explicit outcome model and inverse probability weighting (IPW) simultaneously, with at least one needing to be correct for the estimator to be consistent. While iw uses one period before the treatment’s onset as the reference period, csdid allows users to choose one or multiple pretreatment periods as the reference.We will only implement $\hat{\delta}^{CS-dr}_{ny}$, as most models in our replication sample do not include additional covariates, rendering $\hat{\delta}^{CS-dr}_{nev}$ and $\hat{\delta}^{IW}$ numerically identical when there are no missing data.
@@
The “stacked” DID or regression is another related estimator sometimes used to address HTE concerns. As described by Baker, Larcker and Wang (2022), it involves creating separate sub-datasets for each treated cohort by combining data from that cohort (around treatment adoption) and never-treated cohort data from the same periods. These cohort-specific datasets are then “stacked” to form a single dataset. An event study regression akin to Equation (3) is then run, using sub-dataset specific unit and time dummies. This method is akin to iw and the never-treated version of csdid without covariates; they all use the same data. However, stacked regression estimates a single DTE for a given relative period, rather than separate estimates for each cohort. Essentially, stacked DID is a special case of iw that uses immutable weights selected by OLS. We denote the corresponding estimand $\delta_{l}^{vw}$ to reflect the fact that it is variance-weighted. These weights will be convex, and so stacked DID avoids the “negative weighting” problem, but in general are not proportional to cohort sizes, and thus the stacked DID estimand is not the same as the ATT. This, however, is arguably a minor issue related to external validity, as the estimand will apply to a population with a different distribution over individual treatment effects (ITE).
@@
In settings with treatment reversals, separate groups of researchers have converged on the same strategy for choosing a comparison group: matching treated and control observations that belong to units with identical treatment histories. IKW (2023) suggest one such estimator, PanelMatch, which begins by constructing a “matched set” for each pair $(i,t)$ such that unit $i$ transitions into treatment at time $t$. This matched set includes units that are both (1) not under treatment at time $t$ and (2) share the same treatment history as $i$ for a fixed number of periods leading up to the treatment onset. For each post-period $l$ periods since treatment adoption, it then estimates a local DID. It uses a fixed pre-period $s<t$ and the post-period $t-1+l$. The treatment “group” comprises solely of $(i,t)$, and the members of the matched set for $(i,t)$ that are still under control during that post-period serves as the comparison group. To obtain $\delta_{l}$ for a given $l$, it then averages over the relevant local estimates from all pairs $(i,t)$ such that $(i,t+l-1)$ is still under treatment (i.e., no treatment reversal has occurred yet). IKW (2023) propose incorporating covariates by “refining” matched sets and use $l=0$ as the reference period.
[…]
@@
All DID extensions are built using local, $2\times 2$ DIDs, and their assumptions reflect this. Specifically, they each rely on a form of the PTA, that is, the expected changes in non-treated potential outcomes from one period to the next are equal between the treated and the chosen comparison groups. In Table 1, we refer to all these assumptions collectively as the PTA. We defer readers to the SM for a fuller account of each method’s assumptions.
@@
Imputation estimators do not explicitly estimate local DIDs. Instead, they take the difference between the observed outcome and an imputed counterfactual outcome for each treated observation. The connection to the TWFE model is in the functional form assumption used to impute counterfactual outcomes. Specifically, an imputation estimator first fits a parametric model for the potential outcome under control $Y_{i,t}(0)$—in our case, this is Model (1)—using only control observations $\{(i,t):D_{i,t}=0\}$. It is also through this outcome model that one can adjust for covariates. Then, it imputes $Y_{i,t}(0)$ for all treated observations $\{(i,t):D_{i,t}=1\}$ using the estimated parameters. Finally, it estimates the ITE, $\tau_{i,t}$, for each treated observation $(i,t)$ by calculating the difference between the observation’s observed outcome $Y_{i,t}=Y_{i,t}(1)$ and its imputed counterfactual outcome $\hat{Y}_{i,t}(0)$. Inference for the estimated $\hat{\tau}_{i,t}$ is possible, although uncertainty estimates need to be adjusted to account for the presence of idiosyncratic errors (e.g., Bai and Ng, 2021a). BJS (2023) and LWX (2022) each propose estimators in this category. Each paper proposes a more general framework that nests many models, including TWFE. The latter also introduces several specific imputation estimators, one of which uses the TWFE model, and the authors refer to the resulting estimator as the fixed effect counterfactual estimator, or FEct. [^13] BJS (2023) and LWX (2022) estimate pretreatment DTE in event-study plots differently. The former uses the first period as reference, while the latter uses the average of all pretreatment periods. We prefer the latter method in DTE estimation because the former may produce event-study plots that are visually counterintutitive (Roth, 2024).
@@
Compared to DID extensions, which typically use a single pre-period and, with the exception of csdid, only a subset of units under control at both the pre- and post-periods as a comparison group, imputation estimators use all available control observations to estimate treated counterfactuals. As such, for each unit, the reference period can be understood as (the average of) all pretreatment periods. Intuitively, this approach should result in higher efficiency. In fact, BJS (2023) demonstrate that imputation estimators are the most efficient among all estimators under the condition of homoskedastic errors. Also in contrast to DID extensions, imputation estimators do not directly assume the PTA. Instead, they restrict the expectation of the error terms from a parametric TWFE model. In Table 2, this is denoted as “zero conditional mean” for BSJ (2023) or “strict exogeneity” for LWX (2022). Again, we refer readers to the SM for the formal statements of these assumptions. Collectively, these assumptions imply a form of conditional, baseline randomization of treatment assignment, which in turn implies the PTA (e.g., Blackwell and Glynn, 2018).
@@
Although DID extensions and imputation methods rely on slightly different identification assumptions, such as the PTA and specific constraints on the error terms, they usually lead to similar observable implications. Researchers commonly use the presence or absence of pretrends to judge the plausibility of the PTA. In the classic two-group, two-period setting, if there are data from additional pretreatment periods, researchers can plot the time series of (average) outcomes of each group and visually inspect whether they indeed trend together. The intuition is that if the PTA holds and the outcome trends of the treated and control groups are indeed parallel in pretreatment periods when $Y(0)$’s are observed for all units, then it is plausible that the PTA also holds in the post-treatment periods, when $Y(0)$’s are no longer observable for units in the treatment group. Conversely, differential trends in the pretreatment periods should make us suspicious of the PTA. In more complex settings or where we wish to control for observed confounders, researchers often use estimates of the dynamic effects before and after the onset of treatment, $\delta_{l}$, to construct so-called “event-study plots” to judge the presence of pretrends. If the PTA holds, then pretreatment DTE estimates should be around zero. We provide a more thorough discussion and examples of the event-study plots in the next section when we introduce our procedure.
@@ 4.  Data and Procedure
Next, we assess the robustness of empirical findings from causal panel analyses in political science and compare results obtained using the different methods we have discussed. We will explain our sample selection rules, describe standard practices in the field, and outline our reanalysis approach. Readers can find a more detailed explanation of our sample selection criteria and replication and reanalysis procedure in Sections A.3. and A.4. of the SM.
@@
Our replication sample comprises articles from three leading political science journals—the APSR, AJPS, and JOP—published over a recent six-year span from 2017 to 2022. We initially include all studies, including both long and short articles, that employ panel data analyses with a binary treatment as a crucial component of their causal argument, resulting in a total of 90 articles. After a careful review of each of these articles, we find that 52 articles use a TWFE model similar to Model (1). We then attempt to replicate the main results of these 52 articles and are successful in 37 cases (71.2%). Though a significant proportion of papers failed to replicate, we note that the success rate is still higher than that of Hainmueller, Mummolo and Xu (2019) at 55% and is comparable to that of Lal et al. (2024), which stands at 67%. We credit this to the new replicability standards set by journals. A detailed explanation of how we select the “main model” is provided in Section A.3. of the SM. Table 2 depicts the distribution of successful replications, along with reasons for replication failures, across the various journals.
@@
Table 3 presents an overview of the standard practices and settings in the articles that we successfully replicated. The vast majority of studies in our sample (92%) use DID designs to justify the use of the TWFE model, while the remaining studies advocate for the model’s ability to exploit “within” variations in the data. Out of the 37 articles, five (13%) employ a classic DID design, which includes two-group, two-period designs (three articles) and multi-period DID designs (two articles). 10 articles (27%) use a staggered (but not classic) DID design, while the remaining 22 articles (59.5%) fall into the “general” category, meaning they allow for treatment reversals. Except for four articles, all studies have a continuous outcome of interest. Most studies adopt cluster-robust SEs or panel-corrected SEs (Beck and Katz, 1995), while five articles apply bootstrap procedures for estimating uncertainties. A subset of authors explore alternative model specifications by adding LDVs (six articles), ULTs (11 articles) and higher-than-unit-level time trends (5 articles). Notably, 22 studies use some type of plot—either average outcomes over time, event-study plots, or both—to evaluate the plausibility of the PTA.
@@
Continuous 33 89.2% event-study plots 18 48.6%
@@
To begin, we aim to understand the research setting and data structure. We visualize the patterns of treatment and outcome variables using plots, which are shown in the SM. In this application, treatment reversals clearly take place. Some data is missing (due to redistricting), but the issue does not seem to be severe. We record important details such as the number of observations, units, and time periods, the type of variance estimator, and other specifics of the main model (not shown here). Next, we replicate the main finding, employing both the original variance estimator and a cluster-bootstrap procedure. We also use a bootstrap refinement procedure recommended by Cameron, Gelbach and Miller (2008) to conduct statistical testing; the results are broadly consistent with the ones based on cluster-bootstrapped SEs and CIs.
We then re-estimate the ATT and DTEs using estimators discussed in Section 3. For staggered adoption treatment cases, we apply six estimators: TWFE (with always treated units removed for easier comparisons with other estimators), PanelMatch, the imputation estimator FEct, Stacked DID, iw, and csdid (not-yet-treated). For applications with treatment reversals like Grumbach and Sahn (2020), we use the first three estimators only. The comparison between the TWFE estimate and the other estimates sheds light on whether original findings are sensitive to the presence of HTE. Figures 2(a)-(d) show the results from this example. The similarity between estimates for ATTs in (a) suggests that the original finding is robust to the choice of estimators. The event-study plots from HTE-robust estimators in (c)-(d) are broadly consistent with the event-study plot from TWFE in (b).
@@
Note: Reanalysis of data from Grumbach and Sahn (2020). Subfigures (a): Treatment effect or ATT estimates from multiple methods; subfigures (b)-(d): event-study plots using TWFE, PanelMatch, and FEct; Subfigure (e)-(f): results from the placebo test and test for carryover effects using FEct—the blue points in (g) and red points in (h) represent the holdout periods in the respective tests. CIs for TWFE, PanelMatch, and FEct In all subfigures are produced by bootstrap percentile methods. The robust confident set for the ATT ($\bar{M}=0.5$) is shown in pink in (e).
Next, we conduct diagnostic tests based on FEct to further assess how plausible the PTA is (including the $F$ test and the placebo test) and, in applications with treatment reversal, the no-carryover-effect assumption. We use FEct because it is applicable across all studies in our replication sample and does not discard data. Figures 2(d), (e) and (f) show the results from the $F$ test, placebo test, and test for no carryover effects on our running example, respectively. Both a visual inspection and the formal tests suggest that the PTA and the no-carryover-effect assumption are quite plausible.